OpenStudy (anonymous):

A solid E lies within the cylinder x^2+y^2=1, below the plane z=4, and above the paraboloid z=2-x^2-y^2. The density of the solid is p(x,y,z)=sqrt(x^2+y^2). Calculate the mass of the solid.

OpenStudy (anonymous):

hmmm, debating between a triple integral and a double integral

OpenStudy (anonymous):

i'm going with the triple integral since it's finding a mass

OpenStudy (anonymous):

you have a graphing software?

OpenStudy (anonymous):

nope. I have the answer just not the work behind it. Im taking myself through some triple integral reviews to see if i can figure out how to do it.

OpenStudy (anonymous):

let me graph this thing first

OpenStudy (anonymous):

oh, this is fairly simple

OpenStudy (anonymous):

it will be tripleintegral(sqrt(x^2+y^2))dzdydx the limits for dz is from z=2-x^2-y^2 to z=4 the limits for dy is from y=-sqrt(1-x^2) to y=sqrt(1-x^2) the limit for dx is from x=-1 to x=1

OpenStudy (anonymous):

this is the picture of z=2-x^2+y^2

OpenStudy (anonymous):

That makes more sense. I think i should do some more examples before my final though lol.